OpenStudy (anonymous):

Together, Bob and Tom take 2 hours to get the yard work done. If Bob works alone, it takes him 6 hours. How long would it take Tom to do all of the yard work, working alone? A. 3 hours B. 4 hours C. 5 hours D. 6 hours

OpenStudy (kropot72):

If Bob works alone for one hour, he will get 1/6 of the yard work done. Working alone for two hours, Bob will get 2/6 = 1/3 of the yard work done. When Bob and Tom work together for two hours to finish the yard work, Bob will have done 1/3 of the work, which means Tom must have done 2/3 of the work in the two hours. Therefore if Tom works alone, the time he would take to finish the yard work is given by: \[\large \frac{2}{\frac{2}{3}}=(you\ can\ calculate)\ hours\]